OpenStudy (babyslapmafro):

Please help me find the sum for the following geometric series. (click to see)

OpenStudy (babyslapmafro):

\[\sum_{k=1}^{\infty} \frac{ 4^{k+2} }{ 7^{k-1}}\]

OpenStudy (amistre64):

spose we have a geometric series \[G=1+r+r^2+r^3+...+r^n\] a simple solution to this is to subtract rG from G \[~~~~~~G=1+r+r^2+r^3+...+r^n\\~~~~-rG=-r-r^2-r^3-...-r^n-r^{n+1}\]----------------------------------------\[(1-r)G=1+0+0+0+....+0-r^{n+1}\] therefore\[G=\frac{1-r^{n+1}}{1-r}\] ideally this setup has the top ending at n-1, so thenumerator is 1-r^n

OpenStudy (amistre64):

lets pull out the constant, so that we only have the ^k parts left

OpenStudy (amistre64):

\[\sum_{k=1}^{\infty} \frac{ 4^{k+2} }{ 7^{k-1}}\] \[\sum_{k=1}^{\infty} \frac{ 4^{k} ~4^2 }{ 7^{k}~ 7^{-1}}\] \[\sum_{k=1}^{\infty} (\frac{4}{7})^k~16*7\]

OpenStudy (amistre64):

notice that the 112 pulls out and we are left with G = r+r^2+r^3+r^4+....+r^k -rG = -r^2- .................-r^k-r^(k+1) -------------------------------------- (1-r)G = r - r^(k-1) G = \(\Large \frac{r - r^{(k-1)}}{1-r}\) the limit as k to inf G = \(\Large \frac{r - r^{inf}}{1-r}\)will depend on if 4/7 < 1

OpenStudy (amistre64):

if so, then r^inf = 0, if |r| < 1

OpenStudy (babyslapmafro):

ok thank you for the help i can find the sum from here

OpenStudy (amistre64):

:) have fun